Liquid octane (CH3(CH2)6CH3) will react with gaseous oxygen (O2) to produce gaseous Carbon dioxide (CO2) and gaseous water (H2O). Suppose 10. g of octane is mixed with 69.1g of oxygen. Calculate the minimum mass of octane that could be left over by the chemical reaction


1
Expert's answer
2022-02-28T11:07:12-0500

2C8H18 + 25O2 = 16CO2 + 18H2O

According to the equation, n (O2) = 25 / 2 x n (C8H18).

n = m / M

M (C8H18) = 114.2 g/mol

M (O2) = 32 g/mol


The available amounts of reactants

n (C8H18) = 10 / 114.2 = 0.088 mol

n (O2) = 69.1 / 32 = 2.16 mol


C8H18 is the limiting reactant. The amount of O2 left:

n (O2)left = 2.16 - (25/2 x 0.088) = 1.06 mol

m (O2)left = n x M = 1.06 x 32 = 34 g